# : Timing Interruption Monitoring and Security Assessment for Fronthaul Environments

### v1

@@
Abstract
5G and beyond cellular systems embrace the disaggregation of RAN components, exemplified by the evolution of the fronthual (FH) connection between cellular baseband and radio unit equipment. Crucially, synchronization over the FH is pivotal for reliable 5G services. In recent years, there has been a push to move these links to an Ethernet-based packet network topology, leveraging existing standards and ongoing research for Time-Sensitive Networking (TSN). However, TSN standards, such as Precision Time Protocol (PTP), focus on performance with little to no concern for security. This increases the exposure of the open FH to security risks. Attacks targeting synchronization mechanisms pose significant threats, potentially disrupting 5G networks and impairing connectivity.
@@ Abstract
1 Introduction
@@
Recent advancements in 5G and beyond have focused on disaggregating the Next Generation Node Base (gNB) and core networks, leading to the evolution of more open and modular systems. As a result, the fronthual (FH) has emerged as a critical component, enabling the separation of radio elements from baseband processing. Disaggregated FH deployments were originally based on the Common Public Radio Interface (CPRI). While CPRI facilitated high-speed, centralized processing of radio signals, it lacked built-in security measures due to its design for single-vendor, co-located deployments [1]. The subsequent enhanced CPRI (eCPRI) introduced some security improvements, but significant gaps remain, particularly in synchronization [2]. The eCPRI-based O-RAN ALLIANCE’s open FH now connects disaggregated components from multiple vendors, often using switched Ethernet topologies [3, 4]. While Ethernet’s flexibility supports diverse traffic types, it also exposes the FH to threats, as demonstrated by potential attacks on interconnected devices [5, 6, 7, 8, 9]. These attacks can disrupt synchronization and cause outages. For example, as shown in Fig. 1, an attacker that gains access to the FH network can carry out spoofing attacks that cause synchronization drift, resulting in a reduction in throughput and eventually a complete crash of the base station. Although there are ongoing efforts by industry, government, and academia to address these security gaps [1, 10, 11, 12, 6, 13], no comprehensive standards for securing the FH have yet been established.
@@
The open FH relies on precise timing through the Synchronization Plane (S-plane) to perform critical operations like OFDM synchronization, carrier aggregation, and handovers, which require time accuracy within tight margins [14, 15]. This synchronization depends on the Precision Time Protocol (PTP), which, like the FH, was initially developed without any security mechanisms [16]. While recent updates to the PTP standard include additional security features [17, Annex P], these have not been widely implemented in Open Radio Access Network (Open RAN) environments, leaving vulnerabilities that attackers could exploit.
@@
Motivation and Scope. Security is critical for the success and widespread adoption of any system, and Open RAN has been criticized for lacking robust security mechanisms. In response, the O-RAN ALLIANCE established WG11 [11] to define a security framework, including zero-trust architectures [18], procedures, and defense mechanisms essential for Open RAN’s success [13]. While WG11 acknowledges that attacks on the S-plane can have a high impact, they consider the likelihood low due to the perceived sophistication required by an attacker [19]. However, a significant gap remains in analyzing security risks and proposing solutions for the open FH S-plane in production-grade Open RAN environments. Many researchers lack access to realistic over-the-air systems for testing and validating both attacks and countermeasures. To address this, we developed a Digital Twin system that replicates our production-ready FH network environment, enabling rapid experimentation and analysis.
@@
Experimental Analysis: We conduct the first experimental analysis of the impact of timing attacks in a production-ready private cellular network (see Table II), following O-RAN ALLIANCE procedures [20]. Our findings reveal that these attacks can lead to catastrophic outages of the gNB, highlighting critical vulnerabilities in PTP synchronization mechanisms.
@@
Open Source and Dataset Contributions: Upon acceptance, we will open source our framework, including our Digital Twin design, the attacks, ML models, automation scripts, and a .pcap traffic trace dataset, to support further research and advancement in O-RAN security.
@@
The rest of the paper is organized as follows. In Section 2, we provide background on the S-plane and PTP, while Section 3 surveys prior work. Next, we describe the threat model in Section 4 and our test beds in Section 5. We show the impact of the attacks in Section 6. Then, we demonstrate the need for ML based PTP monitoring tool in Section 7, describe our model in 8 and show state of the art results in Section 9. Finally, we provide additional insights and directions for research in Section 10, and draw our conclusions in Section 11.
2 Background
Traditional Radio Access Network (RAN) architectures are monolithic, leading to vendor lock-in and limited flexibility. The Open RAN paradigm addresses these issues with disaggregated, software-based components and open interfaces [21], based on the 3GPP 7.2 split between Radio Unit (RU) and DU. Current O-RAN/ETSI standards [3, 4] do not mandate encryption on the FH interface due to bandwidth and latency concerns [13], though some studies suggest using Media Access Control Security (MACsec) for added security [5, 22, 23]. We start by defining key acronyms in Table I, followed by background on the S-plane, PTP, and security.
@@ 2 Background
2.1 Fronthaul S-plane
The FH S-plane standards define four clock model and synchronization topologies based on the Lower Layer Split Control Plane (LLS-C): LLS-C1, LLS-C2, LLS-C3, and LLS-C4 [3], as illustrated in Fig. 2. LLS-C1, where the DU and RU are directly connected without a switched network, poses the least security risk due to the absence of intermediate attack vectors. LLS-C4 provides timing directly to the DU and RU using local GNSS receivers, reducing PTP attack opportunities, except for potential GNSS jamming.
In contrast, LLS-C2 and LLS-C3 are more vulnerable due to their reliance on the switched network for PTP message transport. In these scenarios, every device in the switched network represents a potential attack vector. In LLS-C2, the DU is part of the synchronization chain to the RU with one or more Ethernet switches in between. LLS-C3 differs in that the DU is not in the synchronization chain; instead, timing is distributed from the Primary Reference Time Clocks (PRTC) to the DU and RU (s) through the network. The increased complexity and multiple entities involved in securing the network infrastructure expand the risk surface, making LLS-C2 and LLS-C3 the most susceptible to attacks [24, 19].
2.2 PTP Overview
The PTP standard (IEEE 1588 [25]) ensures precise network clock synchronization in the nanosecond range. It achieves this through hardware time stamping, which minimizes delays from the networking stack [26]. The process involves electing a master clock that synchronizes all slave clocks in the network through a series of exchanges and messages.
@@ 2.2 PTP Overview
- ClockClass: Indicates the clock’s traceability and suitability as a time source.
- ClockAccuracy: Reflects the clock’s precision.
@@
This process is repeated periodically and new timestamps ( and ) are taken again as shown in Fig. 3. After receiving two Sync messages, the Slave can calculate the frequency difference relative to the Master clock, which is called . By calculating this drift, the Slave can adjust its own frequency to align with the Master clock, ensuring precise synchronization. To meet all FH timing requirements, additional Physical Layer Frequency Support (PLFS) support, such as SyncE, is required.
Delay refers to the time it takes for a message or signal to travel from one clock to another through the network. In general, delay can be computed via two mechanisms: (i) the E2E mechanism which measures the delay from the master to the slave; and (ii) the P2P mechanism which measures the delay between any two nodes irrespective of their role. E2E does not require intermediate nodes to be PTP aware, while P2P requires all intermediate switches to be PTP aware. In our experiments, we use the E2E mechanism, which is the default configuration. In this case, the master sends a Sync (and optional FollowUp) message and takes the timestamp (). The slave receives the Sync message, takes the timestamp (), sends a DelayReq and takes the timestamp (). The master receives the DelayReq message and takes the timestamp () and stores it. Lastly, the master sends a DelayResp message with the timestamp (). After receiving the DelayResp message from the master, the slave can measure the delay between the two nodes: . This process, also shown in Fig. 3, is repeated at defined intervals.
@@
2.3 PTP Security
The original PTP standard, first published over twenty years ago, did not include any security mechanisms [16]. Subsequent revisions added a security annex, updated in 2019, detailing a flexible, multipronged security approach [25].
@@ 2.3 PTP Security
Prong B addresses external security mechanisms, including MACsec (Layer 2) and IPsec (Layer 3). However, O-RAN Alliance specifications [3] and ITU-T G.8275.1 [27] generally do not permit IP transport. Using MACsec poses challenges; software implementations create virtual interfaces that disrupt hardware timestamping, reducing accuracy significantly. Although hardware-based MACsec solutions have been proposed [28, 29], they increase costs, delay, and complexity, reducing the virtualization central to Open RAN.
Prong C focuses on redundancy (e.g., multiple grandmasters, alternate paths) to mitigate attacks and enhance system resilience. While beneficial, these methods add cost and complexity, conflicting with the push for increased virtualization and decreased costs in Open RAN [21].
Prong D emphasizes monitoring and management. Monitoring PTP performance can help detect potential security attacks [25]. The annex suggests parameters to monitor, but does not specify actions to take upon detecting threats. Given the costs associated with the other prongs, it is logical to implement these security measures only when an attack is detected.
3 Related Work
@@
| [30] | Data center | Spoofing, Comp. Master Clock | None | Experimental testbed |
| [31] | None | Spoofing, Comp. Master Clock | Prong A, Modifications to PTP | Experimental testbed |
| [32] | None | None | Prong A | Experimental testbed |
| [26] | None | None | Prong A | Experimental testbed |
| [33, 34, 35] | Smart Grid | Spoofing, Replay | Modifications to PTP | Formal model verification |
| [16] | IoT | Spoofing | Prong C, Modifications to PTP | IoT testbed |
| [36] | None | Spoofing | Prong C, Modifications to PTP | Experimental testbed |
| [15] | O-RAN | None | None | None |
| [5, 29, 28, 22] | O-RAN | None | Prong B | Experimental testbed |
@@
There is a growing awareness of the need to secure PTP. Several attacks are demonstrated in a data center context by DeCusatis et al. [30]. Itkin and Wool developed a detailed analysis of threats against PTP and proposed using an efficient elliptic-curve public-key signature for Prong A [31]. Shereen et al. [32] experimentally evaluated implementing Prong A, finding that software-based authentication adds at least an additional 70s of delay. Similarly, Rezabek et al. [26] evaluated software-based authentication and concluded that there is visible degradation of clock synchronization for each hop in the network with standard deviations between 118 and 571ns.
There are several works that investigate security threats and propose related defense mechanisms for PTP in smart grids. Moussa et al. [33] proposed adding a new type of PTP clock and modifying the PTP slave functionality. Then in subsequent work, Moussa et al. proposed adding further message types to aid in detection and mitigation [35], and adding feedback from both slaves and master clocks to a reference entity [34].
A method targeting Prong C, where clock and network path redundancy are added, is proposed by Shi et al. [16]. Likewise, the approach by Finkenzeller et al. [36] uses redundant, or cyclic paths, to detect and mitigate time delay attacks.
Maamary et al. give an overview of several threats to the FH S-plane and discuss possible countermeasures [15], though they do not implement either attacks or countermeasures. FH security considerations are further described and MACsec is proposed as a solution by Dik et al. [5, 29, 28, 22]. While these works are the most similar to ours, they primarily address Prong B and could be implemented in parallel to our work.
@@ 4 Threat model
The attack scenario we consider for both the production-ready network and digital twin environment is shown in Fig. 1. In this setup, a switched network connects a software-based DU, with robust computational and storage capabilities (e.g., an edge server running an open-source network stack), to a more limited RU, such as an FPGA-based device. The complexity of the Open RAN architecture, with its many interconnected components, heightens security risks[37]. Although the fronthaul network is protected by IEEE 802.1X, adversaries may bypass this security in a variety of ways.
@@
Increased Remote Access. With its open and multi-vendor architecture, O-RAN relies on a diverse range of suppliers, developers, and service providers, each with varying levels of access and control over networked components. As a result, the number of personnel who can potentially access sensitive infrastructure grows, increasing the risk of misconfigured or compromised authentication processes. IBM’s X-Force Threat Intelligence 2023 report identified misuse of valid credentials (36%) and plaintext credentials on endpoints (33%) as common initial access vectors in cloud environments [38]. Such vulnerabilities are particularly concerning in O-RAN, where distributed network functions and the involvement of multiple parties create more opportunities for configuration errors and unauthorized access to the DUs [6, 39].
@@
Increased Physical Access. The distributed nature of gNB edge servers, often located in accessible public spaces such as sidewalks, rooftops, or building basements, makes on-site physical access an increasingly viable entry point for attacks on the FH [40]. Physical proximity to fronthaul infrastructure lowers the barrier for potential attackers compared to centralized cloud data centers. Once on-site, an adversary can bypass 802.1X by inserting a rogue device (e.g., a mini PC) into the network to intercept, modify, or replay traffic over already authenticated connections.
@@
Dependency and Supply Chain Vulnerabilities. Software-based network functions in 5G, including those in Open RAN, are susceptible to supply chain attacks, particularly with the use of unvetted and globally sourced components. A recent government report highlighted that supply chain risks from untrusted hardware and software vendors are significant threats, as adversarial actors could embed backdoors in either the DU or RU to manipulate fronthaul traffic without being detected by 802.1X [40]. Recent studies of several O-RAN software stacks (including those provided by the open-source projects in the O-RAN Software Community) have uncovered numerous high-risk dependencies, misconfigurations, and weak security practices, such as inadequate access control, lack of encryption, and poor secret management across multiple components [9, 7, 8, 41, 42].
@@
The combination of increased remote access to DUs and physical access to the FH, along with dependency and supply chain vulnerabilities, makes gaining access to a device connected to the FH network not just plausible but increasingly likely. Once an attacker gains access to a device on the switched network, they can observe, replay, or inject PTP broadcast packets sent through the switch.
4.1 Attack Methods

### v2

@@
Abstract.
5G and beyond cellular systems embrace the disaggregation of RAN components, exemplified by the evolution of the fronthaul (FH) connection between cellular baseband and radio unit equipment. Crucially, synchronization over the FH is pivotal for reliable 5G services. In recent years, there has been a push to move these links to an Ethernet-based packet network topology, leveraging existing standards and ongoing research for Time-Sensitive Networking (TSN). However, TSN standards, such as Precision Time Protocol (PTP), focus on performance with little to no concern for security. This increases the exposure of the open FH to security risks. Attacks targeting synchronization mechanisms pose significant threats, potentially disrupting 5G networks and impairing connectivity.
@@ Abstract.
This is the authorsβ version of the work accepted to
ACM Transactions on Privacy and Security (TOPS), 2025.
The final version will be available at https://doi.org/[insertDOIwhenavailable].
1. Introduction
@@
Recent advancements in 5G and beyond have focused on disaggregating the Next Generation Node Base (gNB) and core networks, leading to the evolution of more open and modular systems. As a result, the fronthaul (FH) has emerged as a critical component, enabling the separation of radio elements from baseband processing. Disaggregated FH deployments were originally based on the Common Public Radio Interface (CPRI). While CPRI facilitated high-speed, centralized processing of radio signals, it lacked built-in security measures due to its design for single-vendor, co-located deploymentsΒ (wong2022security). The subsequent enhanced CPRI (eCPRI) introduced some security improvements, but significant gaps remain, particularly in synchronizationΒ (ecpri). The eCPRI-based O-RAN ALLIANCEβs open FH now connects disaggregated components from multiple vendors, often using switched Ethernet topologiesΒ (OranWG4; etsiFh). While Ethernetβs flexibility supports diverse traffic types, it also exposes the FH to threats, as demonstrated by potential attacks on interconnected devicesΒ (dik2021transport; abdalla2022toward; hung2024security; atalay2023securing; 10.1145/3643833.3656118). These attacks can disrupt synchronization and cause outages. For example, as shown in Fig.Β 1, an attacker that gains access to the FH network can carry out spoofing attacks that cause synchronization drift, resulting in a reduction in throughput and eventually a complete crash of the base station. Although there are ongoing efforts by industry, government, and academia to address these security gapsΒ (wong2022security; ericcson; OranWG11-secreqspec; DoD; abdalla2022toward; groen2024securing), no comprehensive standards for securing the FH have yet been established.
@@
The open FH relies on precise timing through the Synchronization Plane (S-plane) to perform critical operations like OFDM synchronization, carrier aggregation, and handovers, which require time accuracy within tight marginsΒ (municio2023ran; s_plane_survey). This synchronization depends on the Precision Time Protocol (PTP), which, like the FH, was initially developed without any security mechanismsΒ (shi2023ms). While recent updates to the PTP standard include additional security featuresΒ (ptp_standard, Annex P), these have not been widely implemented in Open Radio Access Network (Open RAN) environments, leaving vulnerabilities that attackers could exploit.
@@
Motivation and Scope. Security is critical for the success and widespread adoption of any system, and Open RAN has been criticized for lacking robust security mechanisms. In response, the O-RAN ALLIANCE established WG11Β (OranWG11-secreqspec) to define a security framework, including zero-trust architecturesΒ (zta), procedures, and defense mechanisms essential for Open RANβs successΒ (groen2024securing). While WG11 acknowledges that attacks on the S-plane can have a high impact, they consider the likelihood low due to the perceived sophistication required by an attackerΒ (OranWG11-threatmodel). However, a significant gap remains in analyzing security risks and proposing solutions for the open FH S-plane in production-grade Open RAN environments. Many researchers lack access to realistic over-the-air systems for testing and validating both attacks and countermeasures. To address this, we developed a Digital Twin system that replicates our production-ready FH network environment, enabling rapid experimentation and analysis.
@@
Experimental Analysis: We conduct the first experimental analysis of the impact of timing attacks in a production-ready private cellular network (see TableΒ 2), following O-RAN ALLIANCE proceduresΒ (OranWG11-sectestspec). Our findings reveal that these attacks can lead to catastrophic outages of the gNB, highlighting critical vulnerabilities in PTP synchronization mechanisms.
@@
Open Source and Dataset Contributions: We open source our framework, including our Digital Twin design, the attacks, ML models, automation scripts, and a .pcap traffic trace dataset are all open source, to support further research and advancement in O-RAN securityΒ (git_repo).
@@
The rest of the paper is organized as follows. In SectionΒ 2, we provide background on the S-plane and PTP, while SectionΒ 3 surveys prior work. Next, we describe the threat model in SectionΒ 4 and our test beds in SectionΒ 5. We show the impact of the attacks in SectionΒ 6. Then, we demonstrate the need for ML based PTP monitoring tool in SectionΒ 7, describe our model in 8 and show state of the art results in SectionΒ 9. Finally, we provide additional insights and directions for research in SectionΒ 10, and draw our conclusions in SectionΒ 11.
2. Background
Traditional Radio Access Network (RAN) architectures are monolithic, leading to vendor lock-in and limited flexibility. The Open RAN paradigm addresses these issues with disaggregated, software-based components and open interfacesΒ (polese2023understanding), based on the 3GPP 7.2 split between Radio Unit (RU) and DU. Current O-RAN/ETSI standardsΒ (OranWG4; etsiFh) do not mandate encryption on the FH interface due to bandwidth and latency concernsΒ (groen2024securing), though some studies suggest using Media Access Control Security (MACsec) for added securityΒ (dik2021transport; dik2023open; cho2021secure). We start by defining key acronyms in TableΒ 1, followed by background on the S-plane, PTP, and security.
@@ 2. Background
2.1. Fronthaul S-plane
The FH S-plane standards define four clock models and synchronization topologies based on the Lower Layer Split Control Plane (LLS-C): LLS-C1, LLS-C2, LLS-C3, and LLS-C4Β (OranWG4), as illustrated in Fig.Β 2. Each of these profiles reflects a different method for distributing synchronization between the DU and RU, with varying complexity and exposure to attack.
In LLS-C1, the DU and RU are directly connected, typically via a point-to-point Ethernet link. This straightforward setup uses standard PTP signaling between the two endpoints, often with the DU serving as the master and the RU as the slave. With no intermediate devices involved, there are fewer components to misconfigure or compromise, resulting in a relatively low risk of timing manipulation.
LLS-C4 similarly avoids reliance on the transport network by providing each deviceβboth the DU and one or more RUs βwith a local GNSS receiver. These receivers discipline the internal clocks using GNSS signals and holdover algorithms, reducing the need for any in-band synchronization messages. This setup removes network-based attack vectors but introduces potential vulnerabilities to GNSS jamming and spoofing, which must be mitigated with resilient oscillator design and fallback mechanisms.
In contrast, LLS-C2 and LLS-C3 rely heavily on the transport network to carry synchronization information, making them far more vulnerable to attack. LLS-C2 places one or more Ethernet switches between the DU and RU, forming a chain in which each switch must function as a Boundary Clock (BC) or Transparent Clock (TC). These intermediate devices forward or adjust PTP messages in real time, and their correct behavior is critical for accurate timing. Any misconfiguration, software bug, or compromise of these switches can introduce subtle timing errors or deliberate tamperingβboth of which may be difficult to detect. LLS-C3 is even more exposed: the DU and RU both receive timing from a central Grandmaster, such as a Primary Reference Time Clocks (PRTC), via a shared switched network. Because neither endpoint provides timing to the other, the entire synchronization path runs through multiple switches and routers, each of which becomes a potential point of failure or compromise. Moreover, both C2 and C3 require careful configuration of PTP parameters, including delay mechanisms, priority levels, and clock class settings, and they depend on hardware timestamping and accurate Best Master Clock (BMC) selection algorithms to maintain time alignment. This increased complexity not only makes these profiles harder to secure and debug but also expands the attack surface significantly. In practice, any adversary capable of accessing the synchronization pathβwhether through compromised network hardware or malicious softwareβcan degrade or hijack timing across the system, impacting radio coordination, scheduling, and service integrityΒ (groen2023implementing; OranWG11-threatmodel).
2.2. PTP Overview
The PTP standard (IEEE 1588Β (ptp_standard)) ensures precise network clock synchronization in the nanosecond range. It achieves this through hardware time stamping, which minimizes delays from the networking stackΒ (rezabek2022ptp). The process involves electing a master clock that synchronizes all slave clocks in the network through a series of exchanges and messages.
@@ 2.2. PTP Overview
- ClockClass: Indicates the clockβs traceability and suitability as a time source.
- ClockAccuracy: Reflects the clockβs precision.
@@
This process is repeated periodically and new timestamps ( and ) are taken again as shown in Fig.Β 3. After receiving two Sync messages, the Slave can calculate the frequency difference relative to the Master clock, which is called . By calculating this drift, the Slave can adjust its own frequency to align with the Master clock, ensuring precise synchronization. To meet all FH timing requirements, additional Physical Layer Frequency Support (PLFS) support, such as SyncE, is required. See Appendix (Sec.Β A.2) for more background on PLFS.
Delay refers to the time it takes for a message or signal to travel from one clock to another through the network. In general, delay can be computed via two mechanisms: (i) the E2E mechanism which measures the delay from the master to the slave; and (ii) the P2P mechanism which measures the delay between any two nodes irrespective of their role. E2E does not require intermediate nodes to be PTP aware, while P2P requires all intermediate switches to be PTP aware. In our experiments, we use the E2E mechanism, which is the default configuration. In this case, the master sends a Sync (and optional FollowUp) message and takes the timestamp (). The slave receives the Sync message, takes the timestamp (), sends a DelayReq and takes the timestamp (). The master receives the DelayReq message and takes the timestamp () and stores it. Lastly, the master sends a DelayResp message with the timestamp (). After receiving the DelayResp message from the master, the slave can measure the delay between the two nodes: . This process, also shown in Fig.Β 3, is repeated at defined intervals.
@@
2.3. PTP Security
The original PTP standard, first published over twenty years ago, did not include any security mechanismsΒ (shi2023ms). Subsequent revisions added a security annex, updated in 2019, detailing a flexible, multipronged security approachΒ (ptp_standard).
@@ 2.3. PTP Security
Prong B addresses external security mechanisms, including MACsec (Layer 2) and IPsec (Layer 3). However, O-RAN Alliance specificationsΒ (OranWG4) and ITU-T G.8275.1Β (8275.1) generally do not permit IP transport. Using MACsec poses challenges; software implementations create virtual interfaces that disrupt hardware timestamping, reducing accuracy significantly. Although hardware-based MACsec solutions have been proposedΒ (dik2023macsec; dik2023100), they increase costs, delay, and complexity, reducing the virtualization central to Open RAN.
Prong C focuses on redundancy (e.g., multiple grandmasters, alternate paths) to mitigate attacks and enhance system resilience. While beneficial, these methods add cost and complexity, conflicting with the push for increased virtualization and decreased costs in Open RANΒ (polese2023understanding).
Prong D emphasizes monitoring and management. Monitoring PTP performance can help detect potential security attacksΒ (ptp_standard). The annex suggests parameters to monitor, but does not specify actions to take upon detecting threats. Given the costs associated with the other prongs, it is logical to implement these security measures only when an attack is detected.
3. Related Work
@@
| (decusatis2019impact) | Data center | Spoofing, Comp. Master Clock | None | Experimental testbed |
| (itkin2017security) | None | Spoofing, Comp. Master Clock | Prong A, Modifications to PTP | Experimental testbed |
| (shereen2019next) | None | None | Prong A | Experimental testbed |
| (rezabek2022ptp) | None | None | Prong A | Experimental testbed |
| (moussa2016detection; moussa2018securing; moussa2019extension) | Smart Grid | Spoofing, Replay | Modifications to PTP | Formal model verification |
| (shi2023ms) | IoT | Spoofing | Prong C, Modifications to PTP | IoT testbed |
| (finkenzeller2024ptpsec) | None | Spoofing | Prong C, Modifications to PTP | Experimental testbed |
| (s_plane_survey) | O-RAN | None | None | None |
| (dik2021transport; dik2023100; dik2023macsec; dik2023open) | O-RAN | None | Prong B | Experimental testbed |
@@
There is a growing awareness of the need to secure PTP. Several attacks are demonstrated in a data center context by DeCusatis et al.Β (decusatis2019impact). Itkin and Wool developed a detailed analysis of threats against PTP and proposed using an efficient elliptic-curve public-key signature for Prong AΒ (itkin2017security). Shereen et al.Β (shereen2019next) experimentally evaluated implementing Prong A, finding that software-based authentication adds at least an additional 70s of delay. Similarly, Rezabek et al.Β (rezabek2022ptp) evaluated software-based authentication and concluded that there is visible degradation of clock synchronization for each hop in the network with standard deviations between 118 and 571ns.
There are several works that investigate security threats and propose related defense mechanisms for PTP in smart grids. Moussa et al.Β (moussa2016detection) proposed adding a new type of PTP clock and modifying the PTP slave functionality. Then in subsequent work, Moussa et al.Β proposed adding further message types to aid in detection and mitigationΒ (moussa2019extension), and adding feedback from both slaves and master clocks to a reference entityΒ (moussa2018securing).
A method targeting Prong C, where clock and network path redundancy are added, is proposed by Shi et al.Β (shi2023ms). Likewise, the approach by Finkenzeller et al.Β (finkenzeller2024ptpsec) uses redundant, or cyclic paths, to detect and mitigate time delay attacks.
Maamary et al. give an overview of several threats to the FH S-plane and discuss possible countermeasuresΒ (s_plane_survey), though they do not implement either attacks or countermeasures. FH security considerations are further described and MACsec is proposed as a solution by Dik et al.Β (dik2021transport; dik2023100; dik2023macsec; dik2023open). While these works are the most similar to ours, they primarily address Prong B and could be implemented in parallel to our work.
@@ 4. Threat model
The attack scenario we consider for both the production-ready network and digital twin environment is shown in Fig.Β 1. In this setup, a switched network connects a software-based DU, with robust computational and storage capabilities (e.g., an edge server running an open-source network stack), to a more limited RU, such as an FPGA-based device. The complexity of the Open RAN architecture, with its many interconnected components, heightens security risks(nisReport; nokia). Although the fronthaul network is protected by IEEE 802.1X, adversaries may bypass this security in a variety of ways.
@@
Increased Remote Access. With its open and multi-vendor architecture, O-RAN relies on a diverse range of suppliers, developers, and service providers, each with varying levels of access and control over networked components. As a result, the number of personnel who can potentially access sensitive infrastructure grows, increasing the risk of misconfigured or compromised authentication processesΒ (nokia). IBMβs X-Force Threat Intelligence 2023 report identified misuse of valid credentials (36%) and plaintext credentials on endpoints (33%) as common initial access vectors in cloud environmentsΒ (ibm2023). Such vulnerabilities are particularly concerning in O-RAN, where distributed network functions and the involvement of multiple parties create more opportunities for configuration errors and unauthorized access to the DUsΒ (abdalla2022toward; verizon).
@@
Increased Physical Access. The distributed nature of gNB edge servers, often located in accessible public spaces such as sidewalks, rooftops, or building basements, makes on-site physical access an increasingly viable entry point for attacks on the FHΒ (xing2024criticality). Physical proximity to fronthaul infrastructure lowers the barrier for potential attackers compared to centralized cloud data centers. Once on-site, an adversary can bypass 802.1X by inserting a rogue device (e.g., a mini PC) into the network to intercept, modify, or replay traffic over already authenticated connections.
@@
Dependency and Supply Chain Vulnerabilities. Software-based network functions in 5G, including those in Open RAN, are susceptible to supply chain attacks, particularly with the use of unvetted and globally sourced componentsΒ (nokia). A recent government report highlighted that supply chain risks from untrusted hardware and software vendors are significant threats, as adversarial actors could embed backdoors in either the DU or RU to manipulate fronthaul traffic without being detected by 802.1XΒ (xing2024criticality). Recent studies of several O-RAN software stacks (including those provided by the open-source projects in the O-RAN Software Community) have uncovered numerous high-risk dependencies, misconfigurations, and weak security practices, such as inadequate access control, lack of encryption, and poor secret management across multiple componentsΒ (10.1145/3643833.3656118; hung2024security; atalay2023securing; tiberti2022impact; 10588929).
@@
The combination of increased remote access to DUs and physical access to the FH, along with dependency and supply chain vulnerabilities, makes gaining access to a device connected to the FH network not just plausible but increasingly likelyΒ (nokia). Once an attacker gains access to a device on the switched network, they can observe, replay, or inject PTP broadcast packets sent through the switch.
4.1. Attack Methods